Circular No. 6649

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 E. Costa, M. Feroci, L. Piro, P. Soffitta, L. Amati, and M. N. Cinti,

Istituto Astrofisica Spaziale, CNR, Frascati; F. Frontera and G. Zavattini, Universitá di Ferrara; L. Nicastro and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; A. Tesseri, G. Gandolfi, M. Smith, D. Ricci, and A. Coletta, BeppoSAX, Rome; J. Heise and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht; and M. Tavani, Columbia Astrophysics Laboratory and Istituto di Fisica Cosmica e Tecnologie Relative, CNR, Milano, report: “The BeppoSAX Gamma Ray Burst Monitor (GRBM) was triggered on May 8.904 UT. A quick-look analysis of the GRBM light curve shows a rather weak γ-ray burst (GRB) lasting ∼ 15 s. The intensity at its maximum is ∼ 450 counts/s. The GRB was also detected by the Wide Field Camera unit 2 with a maximum countrate of ∼ 130 counts/s. Taking into account the off-axis position, the peak flux corresponds to ∼ 1 Crab. From the WFC image, we have derived a preliminary position of α = 6^{h}54^{m}21^{s}, δ = +79^{o}15′.8 (equinox 2000.0), with a 5′ error radius. One catalogued source lies within the GRB WFC error box: 1RXS J065213.8+790855 is ∼ 8′ from the GRB error-box center. Observations in all bands are solicited.”

SUPERNOVA 1997cl IN ANONYMOUS GALAXY S. Benetti, I. Hook, P. Shaver, and M. Turatto, European Southern

Observatory (ESO), communicate: “A CCD spectrogram (range 374–691 nm, resolution 1.6 nm), obtained on May 9.09 UT with the ESO 3.6-m telescope (+ EFOSC1) at La Silla, shows that SN 1997cl (cf. IAUC 6647) is a peculiar type-IIn supernova. The spectrum is dominated by a relatively strong Hα emission on a relatively red continuum. The Hα profile appears to consist of three components: one unresolved (centered at 687.4 nm), one intermediate (FWHM 1500 km/s), and possibly a broad one (FWHM ∼ 10 000 km/s). A faint Hβ emission shows only the unresolved component. The Fe-II 516.9- and 501.8-nm lines are also visible with P-Cyg profiles, from which an expansion velocity of ∼ 1900 km/s is derived. The redshift of the parent galaxy (measured close to the supernova position, using of Hα and [N II] 658.4-nm emission lines and Ca II H and K absorption lines) is 0.047.”

1997 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green